Diagonal of a rectangle is 15 m. Longer side of rectangle is 12 m. What is perimeter of the rectangle?
Answer by mananth(16946)

Diagonal of a rectangle is 15 m. Longer side of rectangle is 12 m. What is perimeter of the rectangle?
The diagonal and the two adjacent sides form a right triangle
Applying the pytahagoras theorem
Diagonal^2 = Length^2 + width^2
15^2= 12^2 +w^2
15^2-12^2 = width^2
225-144 = width^2
81=width^2
9= width
9*2 + 12*2 = perimeter
18+24 = 42 meters